A racecar starts from rest and travels east along a straight and level track. For the first 5.0 s of the car’s motion, the eastward component of the car’s velocity is given by 𝑣(𝑑) = (10.860π‘š/𝑠 3 )𝑑 2 . What is the acceleration of the car when 𝑣π‘₯ = 12.0π‘š/𝑠?


1
Expert's answer
2021-09-25T13:10:49-0400

Given:

"v(t)=10.860t^2\:\rm m/s"


The acceleration

"a(t)=v'(t)=21.72t"

The velocity of racecar is equal "v=12.0\:\rm m/s" for the time "t=\sqrt{12.0/10.86}=1.051\:\rm s". At this instant it acceleration

"a=21.72*1.051=22.83\:\rm m/s^2"